Circular No. 7615

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001bc, 2001bd, 2001be E. Cappellaro, G. Altavilla, S. Benetti, A. Pastorello, and M. Turatto,

Osservatorio di Padova; A. Clocchiatti, Pontificia Universidad Catolica de Chile; I. J. Danziger and P. Mazzali, Osservatorio di Trieste; and F. Patat, European Southern Observatory (ESO), report the discovery of three new supernovae on CCD frames taken during a search performed with the ESO/MPI 2.2-m telescope (+ Wide Field Imager). The supernovae were identified by subtracting reference images taken with the same equipment during 1999.

SN  2001 UT  α_{2000}  δ_{2000}  Mag.  z Type
2001bc Apr. 20.7 9^{h}46^{m}07.^{s}06 −21^{o}11′28′′.6 23.4 B 0.195 II
2001bd Apr. 19.3 19 13 10.97 −64 17 07.0 22.6 V  0.096 II
2001be Apr. 20.2 9 52 51.21 −20 34 00.0 21.1 V  0.241 Ia

SN 2001bc and SN 2001be are each within 1′′ of the nucleus of its parent galaxy. SN 2001bd lies 3′′.1 west and 1′′.8 south of its galaxy’s nucleus. Each object’s classification and redshift were derived from spectra obtained on Apr. 28 with the European Southern Observatory 8.2-m ANTU Very Large Telescope (+ FORS1).

SUPERNOVA 2001az IN UGC 10483 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001az (cf. IAUC 7614), obtained by P. Berlind on Apr. 30.39 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova before maximum light. Narrow Hα emission from a separate exposure of the nucleus of UGC 10483 yields a recession velocity of 12 200 km/s for the host galaxy. Using this value, the expansion velocity of the supernova is ∼ 12 400 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 4 ± 2 days before maximum light.

R. Chornock, M. Modjaz, and A. V. Filippenko, University of California at Berkeley, report that inspection of a CCD spectrum (range 320–1000 nm), obtained on Apr. 30 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001az is of type Ia, within one week of maximum brightness.

2001 May 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green